# Implicit Safe Set Algorithm for Provably Safe Reinforcement Learning

### v1

@@ 7 Theoretical Results for ISSA When the Sampling Time is Non-Negligible
If the control system satisfies the assumptions in Assumption 1 and Assumption 3, and the safety index design follows the rule described in Section 6.1, the discrete-time system implicit safe set algorithm described in Section 6.2 guarantees the forward invariance and finite time convergence to the set $\mathcal{S}\subseteq\mathcal{X}_{S}$.
@@
7.1 Feasibility of Safety Index for Discrete-Time System
@@
If the dynamic system satisfies the assumptions in and Assumption 1 and Assumption 3, then the discrete-time safety index design rule in Section 6.1 ensures that the robot system in 2D plane has nonempty set of safe control at any state, i.e., $\mathcal{U}_{S}^{D}(x)\neq\emptyset,\forall x$.
@@
$\forall x,\exists u,\text{s.t. }\phi(f(x,u))\leq\phi(x)-\eta~{},$ (29)
@@
$\phi(f(x,u))=\phi(x)+dt\cdot\dot{\phi}(x,u)+dt^{2}\cdot\frac{\ddot{\phi}(x,u)}{2}+\Delta,$ (30)
@@ 7.1.1 Preliminary Results
If the dynamic system satisfies the assumptions in Assumption 1 and Assumption 3, and there’s only one obstacle in the environment, then the safety index design rule for discrete-time systems in Section 6.1 ensures the approximation error $\Delta$ for discrete-time systems is upper bonded by $\Delta_{max}dt^{2}(e^{|w|dt}-1)$, where $\Delta_{max}$ is a positive constant, $dt$ is the discrete time step, and $w$ is the applied angular velocity.
@@
First, we will discuss the derivatives of $d$, which is the relative distance between the robot and the obstacle. According to the zero-order hold assumption for acceleration and angular velocity from Assumption 3, we have $\dot{a}=0$, and $\dot{w}=0$. And we already have $\dot{d}=-v\cos(\alpha)$, then following condition holds:
$d^{(i)}=(i-1)a(\pm\cos(\alpha)|\pm\sin(\alpha))w^{i-2}+v(\pm\sin(\alpha)|\pm\cos(\alpha))w^{i-1}$ (31)
where ±$\pm\cos(\alpha)|\pm\sin(\alpha)$ denotes one of the terms from set
@@
$\phi_{t+1}=\phi_{t}+\frac{\dot{\phi}_{t}}{1!}dt+\frac{\ddot{\phi}_{t}}{2!}dt^{2}+\frac{\dddot{\phi}_{t}}{3!}dt^{3}+\cdots$ (34)
@@
$\frac{\phi_{t}^{(i)}}{i!}dt^{i}=dt^{2}\frac{(i-1)a\cos(\alpha)-(v+kia)\sin(\alpha)w-kv\cos(\alpha)w^{2}}{i!}(wdt)^{i-2}$ (35)
@@
$\frac{(i-1)a\cos(\alpha)-(v+kia)\sin(\alpha)w-kv\cos(\alpha)w^{2}}{i!}$ (36)
@@
where $a_{m}=\max\{-a_{min},a_{max}\}$ and $w_{m}=\max\{-w_{min},w_{max}\}$. Let’s define $\Delta_{max}=(a_{m}+v_{max}w_{m})(1+kw_{m})$, and $\Delta_{max}$ is a constant given a specific $k$ value. Based on (36) and (34), the upper bond of $\frac{\phi_{t}^{(i)}}{i!}dt^{i}$ is defined as following:
$\frac{\phi_{t}^{(i)}}{i!}dt^{i}$ (37)
@@
[…]
@@
If the dynamic system satisfies the assumptions in Assumption 1 and Assumption 3, and there’s only one obstacle in the environment, then the safety index design rule for discrete-time systems in Section 6.1 ensures that the robot system in 2D plane has nonempty set of safe control at any state.
@@
$\forall x,\exists u,\text{s.t. }\ddot{d}\geq\frac{\frac{\eta+dt^{2}\cdot\frac{\ddot{\phi}(x,u)}{2}+\Delta}{dt}-nd^{n-1}\dot{d}}{k}~{}.$ (39)
@@
$\forall(\alpha,v),\exists(a,w),$ s.t. (40)
@@
$\forall(\alpha,v),\exists(a,w),$ s.t. (41)
@@
$\forall(\alpha,v),\exists(a),\text{ s.t. }-a\cos(\alpha)$ $\geq\frac{\frac{\eta}{dt}+dt\cdot\frac{a\cos(\alpha)}{2}+v\cos(\alpha)}{k}$ (42)
@@
$\forall(\alpha,v),\exists(a>0),\text{ s.t. }-a\cos(\alpha)$ $=a|\cos(\alpha)|$ (43)
@@
$\max_{a}-a=-a_{min}\geq\frac{\frac{\eta_{0}}{dt}+v_{\max}}{k}\geq\frac{\frac{\eta_{0}}{dt}+dt\cdot\frac{a}{2}+v_{\max}}{k}\geq\frac{\frac{\eta_{0}}{dt}+dt\cdot\frac{a}{2}+v}{k}$ (44)
@@
$\forall(\alpha,v),\exists(a<0),\text{ s.t. }-a\cos(\alpha)$ $\geq\frac{\frac{\eta_{0}}{dt}+dt\cdot\frac{a}{2}+v}{k}\cos(\alpha)$ (45)
@@
Combine these two cases, we have proved that if the control system satisfies the assumptions in Assumption 1 and Assumption 3, then the safety index design rule in Section 6.1 ensures that the robot system in 2D plane has nonempty set of safe control at any state.
7.1.2 Proof of Proposition 3
@@
Case 1: Firstly, we consider the case where $\phi(x)\geq-\Delta_{\phi_{max}}$. We have $\phi(x)=\max_{i}\phi_{i}(x)\geq-\Delta_{\phi_{max}}$, where $\phi_{i}$ is the safety index with respect to the $i$-th obstacle. According to assumption 3, we have that at any given time, there can at most be one obstacle becoming safety critical, such that $\phi\geq-\Delta_{\phi_{max}}$ (Sparse Obstacle Environment). Therefore, $\max_{i}\phi_{i}(x)>0$ indicates there’s only one obstacle ($j$-th obstacle) in the environment that $\phi_{j}(x)\geq-\Delta_{\phi_{max}}$. Whereas for the rest of the obstacles, we have $\phi_{k}(x)<-\Delta_{\phi_{max}},k\neq j$.
Denote ${\mathcal{U}_{S}^{D}}_{j}(x):=\{u\in\mathcal{U}\mid\phi_{j}(f(x,u))\leq\max\{\phi_{j}(x)-\eta,0\}\}$ According to Lemma 8, we have if the dynamic system satisfies the assumptions in Assumption 3, then the safety index design rule in Section 6.1 ensures ${\mathcal{U}_{S}^{D}}_{j}(x)$ is nonempty. Since $\Delta_{\phi_{max}}$ is the maximum possible $\phi$ change magnitude at one time step, we also have that $\phi_{k}(f(x,u))\leq 0\leq\max\{\phi_{j}(x)-\eta,0\}$, for $u\in{\mathcal{U}_{S}^{D}}_{j}(x)$.
@@
$\phi(f(x,u))=\max_{i}\phi_{i}(f(x,u))\leq 0=\max\{\phi(x)-\eta,0\}$ (47)
In summary, if the dynamic system satisfies the assumptions in Assumption 1 and Assumption 3, then the discrete-time safety index design rule in Section 6.1 ensures that the robot system in 2D plane has nonempty set of safe control at any state, i.e., $\mathcal{U}_{S}^{D}(x)\neq\emptyset,\forall x$.
7.2 Feasibility of Implicit Safe Set Algorithm for Discrete-Time System
@@
If the control system satisfies the assumptions in Assumption 1 and Assumption 3, and there’s only one obstacle in the environment, the safety index design follows the rule described in Section 6.1, the implicit safe set algorithm with convergence trigger algorithm together guarantee the finite time convergence to the set $\mathcal{X}_{S}^{D}$.
@@ 7.2.1 Preliminary Results
If the dynamic system satisfies the assumptions in Assumption 1 and Assumption 3, and there’s only one obstacle in the environment, then the safety index design rule in Section 6.1 ensures that the robot system in 2D plane can always generate safe control such that angular velocity is non-zero for states where $|\sin(\alpha)|\geq\frac{1}{2},v\geq\frac{v_{max}}{2}$.
@@
$\nabla_{|w|}F(0)$ $=(\frac{kv}{dt}+\frac{v+2ka}{2})|\sin(\alpha)|-dt\Delta_{max}$ (50)
@@
According to the safety index design rule in Section 6.1, we have $\frac{v_{\max}}{k}\leq\frac{\frac{\eta_{0}}{dt}+v_{\max}}{k}\leq\min\{-a_{min},a_{max}\}\leq-a_{min}$, which indicates $0<\frac{1}{k}\leq\frac{-a_{min}}{v_{max}}$. We also have $\frac{a_{min}}{2}+\frac{v_{max}}{4dt}>(a_{m}+v_{max}w_{m})(-\frac{a_{min}}{v_{max}}+w_{m})dt$ according to Assumption 3. Therefore, when $v>\frac{v_{max}}{2}$, the lower bound for (50) is summarized as following:
$\nabla_{|w|}F(0)$ $>k(\frac{\frac{v}{dt}+a}{2}-dt(a_{m}+v_{max}w_{m})(\frac{1}{k}+w_{m}))$ (51)
@@ 7.2.2 Proof of Proposition 4
Case 2.2: $|\cos(\alpha)|<\min\{\frac{\sqrt{3}}{2},\frac{\Delta_{\min}}{2}\}$ and $v\geq\frac{v_{max}}{2}$. According to Assumption 3, we have that $|a_{t}|$ is constant between two consecutive time steps, which indicates the following property holds:
@@
$|\cos(\alpha_{t+1})|$ $=|\cos(\alpha_{t})+\Delta_{t}\cos(\alpha)|$ (52)
@@
7.3 Proof of Theorem 2
7.3.1 Preliminary Results
@@
If the control system satisfies the assumptions in Assumption 1 and Assumption 3, and the safety index design follows the rule described in Section 6.1, the implicit safe set algorithm with convergence trigger algorithm together guarantee the forward invariance and finite time convergence to the set $\mathcal{X}_{S}^{D}$.
@@
Firstly, we will prove the forward invariance of set $\mathcal{X}_{S}^{D}$. If the control system satisfies the assumptions in Assumption 1 and Assumption 3, and the safety index design follows the rule described Section 6.1, then we can ensure the system has nonempty set of safe control at any state by Proposition 3. By Proposition 2, implicit safe set algorithm can always find local optima solution of (5). The local optima solution always satisfies the constraint $\phi(f(x_{t},u_{t}))\leq\max\{\phi(x_{t})-\eta,0\}$, which indicates that if $\phi(x_{t_{0}})\leq 0$, then $\phi(x_{t})\leq 0,\forall t\geq t_{0}$ since $\eta\geq 0$. Note that $\phi(x)\leq 0$ demonstrates that $x\in\mathcal{X}_{S}^{D}$. Thus the forward invariance of $\mathcal{X}_{S}^{D}$ is proved.
Secondly, we will prove the finite time convergence to $\mathcal{X}_{S}^{D}$. Suppose at time step $t_{0}$, we have $\phi(x_{t_{0}})=\max_{i}\phi_{i}(x_{t_{0}})>0$. According to Sparse Obstacle Environment assumption from Assumption 3, there’s only one obstacle ($j$-th obstacle) in the environment that $\phi_{j}(x_{t_{0}})>0$. Whereas for the rest of the obstacles, we have $\phi_{k}(x_{t_{0}})<-\Delta_{\phi_{max}},k\neq j$.

### v2

@@ 7. Theoretical Results for ISSA When the Sampling Time is Non-Negligible
If the control system satisfies the assumptions in 1 and 3, and the safety index design follows the rule described in Section 6.1, the discrete-time implicit safe set algorithm described in Section 6.2 guarantees the forward invariance and finite time convergence to the set $\mathcal{S}\subseteq\mathcal{X}_{S}$.
@@
7.1. Feasibility of Safety Index for Discrete-Time System
@@
If the dynamic system satisfies the assumptions in and 1 and 3, then the discrete-time safety index design rule in Section 6.1 ensures that the robot system in 2D plane has nonempty set of safe control at any state, i.e., $\mathcal{U}_{S}^{D}(x)\neq\emptyset,\forall x$.
@@
(29) $\forall x,\exists u,\text{s.t. }\phi(f(x,u))\leq\phi(x)-\eta~,$
@@
(30) $\phi(f(x,u))=\phi(x)+dt\cdot\dot{\phi}(x,u)+dt^{2}\cdot\frac{\ddot{\phi}(x,u)}{2}+\Delta,$
@@ 7.1.1. Preliminary Results
If the dynamic system satisfies the assumptions in 1 and 3, and there’s only one obstacle in the environment, then the safety index design rule for discrete-time systems in Section 6.1 ensures the approximation error $\Delta$ for discrete-time systems is upper bounded by $\Delta_{max}dt^{2}(e^{|w|dt}-1)$, where $\Delta_{max}$ is a positive constant, $dt$ is the discrete time step, and $w$ is the applied angular velocity.
@@
First, we will discuss the derivatives of $d$, which is the relative distance between the robot and the obstacle. According to the zero-order hold assumption for acceleration and angular velocity from 3, we have $\dot{a}=0$, and $\dot{w}=0$. And we already have $\dot{d}=-v\cos(\alpha)$, then following condition holds:
(31) $d^{(i)}=(i-1)a(\pm\cos(\alpha)|\pm\sin(\alpha))w^{i-2}+v(\pm\sin(\alpha)|\pm\cos(\alpha))w^{i-1}$
where $\pm\cos(\alpha)|\pm\sin(\alpha)$ denotes one of the terms from set
@@
(34) $\phi_{t+1}=\phi_{t}+\frac{\dot{\phi}_{t}}{1!}dt+\frac{\ddot{\phi}_{t}}{2!}dt^{2}+\frac{\dddot{\phi}_{t}}{3!}dt^{3}+\cdots$
@@
(35) $\frac{\phi_{t}^{(i)}}{i!}dt^{i}=dt^{2}\frac{(i-1)a\cos(\alpha)-(v+kia)\sin(\alpha)w-kv\cos(\alpha)w^{2}}{i!}(wdt)^{i-2}$
@@
(36) $\frac{(i-1)a\cos(\alpha)-(v+kia)\sin(\alpha)w-kv\cos(\alpha)w^{2}}{i!}$
@@
where $a_{m}=\max\{-a_{min},a_{max}\}$ and $w_{m}=\max\{-w_{min},w_{max}\}$. Let’s define $\Delta_{max}=(a_{m}+v_{max}w_{m})(1+kw_{m})$, and $\Delta_{max}$ is a constant given a specific $k$ value. Based on (36) and (34), the upper bound of $\frac{\phi_{t}^{(i)}}{i!}dt^{i}$ is defined as following:
(37) $\frac{\phi_{t}^{(i)}}{i!}dt^{i}$
@@
Recall that $\phi(x_{t+1})=\phi(x_{t})+dt\cdot\dot{\phi}(x_{t})+dt^{2}\cdot\frac{\ddot{\phi}(x_{t})}{2}+\Delta$, where $\Delta$ is the approximation error due to discrete-time systems, and $\Delta=\frac{\dddot{\phi}_{t}}{3!}dt^{3}+\frac{\ddddot{\phi}_{t}}{4!}dt^{4}+\cdots$. According to the Taylor series of exponential function $e^{x}=1+x+\frac{x^{2}}{2!}+\frac{x^{3}}{3!}+\cdots$, and leverage the result from (37), the upper bound of $\Delta$ can be derived as following:
(38) $\Delta$ $=\frac{\dddot{\phi}_{t}}{3!}dt^{3}+\frac{\ddddot{\phi}_{t}}{4!}dt^{4}+\cdots$
@@
If the dynamic system satisfies the assumptions in 1 and 3, and there’s only one obstacle in the environment, then the safety index design rule for discrete-time systems in Section 6.1 ensures that the robot system in 2D plane has nonempty set of safe control at any state.
@@
(39) $\forall x,\exists u,\text{s.t. }\ddot{d}\geq\frac{\frac{\eta+dt^{2}\cdot\frac{\ddot{\phi}(x,u)}{2}+\Delta}{dt}-nd^{n-1}\dot{d}}{k}~.$
@@
(40) $\forall(\alpha,v),\exists(a,w),$ s.t.
@@
(41) $\forall(\alpha,v),\exists(a,w),$ s.t.
@@
(42) $\forall(\alpha,v),\exists(a),\text{ s.t. }-a\cos(\alpha)$ $\geq\frac{\frac{\eta}{dt}+dt\cdot\frac{a\cos(\alpha)}{2}+v\cos(\alpha)}{k}$
@@
(43) $\forall(\alpha,v),\exists(a>0),\text{ s.t. }-a\cos(\alpha)$ $=a|\cos(\alpha)|$
@@
(44) $\max_{a}-a=-a_{min}\geq\frac{\frac{\eta_{0}}{dt}+v_{\max}}{k}\geq\frac{\frac{\eta_{0}}{dt}+dt\cdot\frac{a}{2}+v_{\max}}{k}\geq\frac{\frac{\eta_{0}}{dt}+dt\cdot\frac{a}{2}+v}{k}$
@@
(45) $\forall(\alpha,v),\exists(a<0),\text{ s.t. }-a\cos(\alpha)$ $\geq\frac{\frac{\eta_{0}}{dt}+dt\cdot\frac{a}{2}+v}{k}\cos(\alpha)$
@@
Combine these two cases, we have proved that if the control system satisfies the assumptions in 1 and 3, then the safety index design rule in Section 6.1 ensures that the robot system in 2D plane has nonempty set of safe control at any state.
7.1.2. Proof of Proposition 2
@@
Case 1: Firstly, we consider the case where $\phi(x)\geq-\Delta_{\phi_{max}}$. We have $\phi(x)=\max_{i}\phi_{i}(x)\geq-\Delta_{\phi_{max}}$, where $\phi_{i}$ is the safety index with respect to the $i$-th obstacle. According to 3, we have that at any given time, there can at most be one obstacle becoming safety critical, such that $\phi\geq-\Delta_{\phi_{max}}$ (Sparse Obstacle Environment). Therefore, $\max_{i}\phi_{i}(x)>0$ indicates there’s only one obstacle ($j$-th obstacle) in the environment that $\phi_{j}(x)\geq-\Delta_{\phi_{max}}$. Whereas for the rest of the obstacles, we have $\phi_{k}(x)<-\Delta_{\phi_{max}},k\neq j$.
Denote ${\mathcal{U}_{S}^{D}}_{j}(x):=\{u\in\mathcal{U}\mid\phi_{j}(f(x,u))\leq\max\{\phi_{j}(x)-\eta,0\}\}$ . According to Lemma 4, we have if the dynamic system satisfies the assumptions in 3, then the safety index design rule in Section 6.1 ensures ${\mathcal{U}_{S}^{D}}_{j}(x)$ is nonempty. Since $\Delta_{\phi_{max}}$ is the maximum possible $\phi$ change magnitude at one time step, we also have that $\phi_{k}(f(x,u))\leq 0\leq\max\{\phi_{j}(x)-\eta,0\}$, for $u\in{\mathcal{U}_{S}^{D}}_{j}(x)$.
@@
(47) $\phi(f(x,u))=\max_{i}\phi_{i}(f(x,u))\leq 0=\max\{\phi(x)-\eta,0\}$
In summary, if the dynamic system satisfies the assumptions in 1 and 3, then the discrete-time safety index design rule in Section 6.1 ensures that the robot system in 2D plane has nonempty set of safe control at any state, i.e., $\mathcal{U}_{S}^{D}(x)\neq\emptyset,\forall x$.
7.2. Feasibility of Implicit Safe Set Algorithm for Discrete-Time System
@@
If the control system satisfies the assumptions in 1 and 3, and there’s only one obstacle in the environment, the safety index design follows the rule described in Section 6.1, the implicit safe set algorithm with convergence trigger algorithm together guarantee the finite time convergence to the set $\mathcal{X}_{S}^{D}$.
@@ 7.2.1. Preliminary Results
If the dynamic system satisfies the assumptions in 1 and 3, and there’s only one obstacle in the environment, then the safety index design rule in Section 6.1 ensures that the robot system in 2D plane can always generate safe control such that angular velocity is non-zero for states where $|\sin(\alpha)|\geq\frac{1}{2},v\geq\frac{v_{max}}{2}$.
@@
(50) $\nabla_{|w|}F(0)$ $=(\frac{kv}{dt}+\frac{v+2ka}{2})|\sin(\alpha)|-dt\Delta_{max}$
@@
According to the safety index design rule in Section 6.1, we have $\frac{v_{\max}}{k}\leq\frac{\frac{\eta_{0}}{dt}+v_{\max}}{k}\leq\min\{-a_{min},a_{max}\}\leq-a_{min}$, which indicates $0<\frac{1}{k}\leq\frac{-a_{min}}{v_{max}}$. We also have $\frac{a_{min}}{2}+\frac{v_{max}}{4dt}>(a_{m}+v_{max}w_{m})(-\frac{a_{min}}{v_{max}}+w_{m})dt$ according to 3. Therefore, when $v>\frac{v_{max}}{2}$, the lower bound for (50) is summarized as following:
(51) $\nabla_{|w|}F(0)$ $>k(\frac{\frac{v}{dt}+a}{2}-dt(a_{m}+v_{max}w_{m})(\frac{1}{k}+w_{m}))$
@@ 7.2.2. Proof of Proposition 5
Case 2.2: $|\cos(\alpha)|<\min\{\frac{\sqrt{3}}{2},\frac{\Delta_{\min}}{2}\}$ and $v\geq\frac{v_{max}}{2}$. According to 3, we have that $|a_{t}|$ is constant between two consecutive time steps, which indicates the following property holds:
@@
(52) $|\cos(\alpha_{t+1})|$ $=|\cos(\alpha_{t})+\Delta_{t}\cos(\alpha)|$
@@
7.3. Proof of Theorem 1
7.3.1. Preliminary Results
@@
If the control system satisfies the assumptions in 1 and 3, and the safety index design follows the rule described in Section 6.1, the implicit safe set algorithm with convergence trigger algorithm together guarantee the forward invariance and finite time convergence to the set $\mathcal{X}_{S}^{D}$.
@@
Firstly, we will prove the forward invariance of set $\mathcal{X}_{S}^{D}$. If the control system satisfies the assumptions in 1 and 3, and the safety index design follows the rule described Section 6.1, then we can ensure the system has nonempty set of safe control at any state by Proposition 2. By Proposition 5, implicit safe set algorithm can always find local optima solution of (5). The local optima solution always satisfies the constraint $\phi(f(x_{t},u_{t}))\leq\max\{\phi(x_{t})-\eta,0\}$, which indicates that if $\phi(x_{t_{0}})\leq 0$, then $\phi(x_{t})\leq 0,\forall t\geq t_{0}$ since $\eta\geq 0$. Note that $\phi(x)\leq 0$ demonstrates that $x\in\mathcal{X}_{S}^{D}$. Thus the forward invariance of $\mathcal{X}_{S}^{D}$ is proved.
Secondly, we will prove the finite time convergence to $\mathcal{X}_{S}^{D}$. Suppose at time step $t_{0}$, we have $\phi(x_{t_{0}})=\max_{i}\phi_{i}(x_{t_{0}})>0$. According to Sparse Obstacle Environment assumption from 3, there’s only one obstacle ($j$-th obstacle) in the environment that $\phi_{j}(x_{t_{0}})>0$. Whereas for the rest of the obstacles, we have $\phi_{k}(x_{t_{0}})<-\Delta_{\phi_{max}},k\neq j$.